# Florence, Italy - Pisa, Italy

### v1

@@ 4.4 Results on Generalization to Unseen Classes
In Table 3 we give results for the unseen class generalization task. In these experiments each dataset is split into 50% of the classes as a base for training few-shot adaptation, and the remaining 50% as new classes on which zero-shot performance is evaluated. KDPL consistently outperforms the corresponding baseline methods for both backbones, demonstrating improvement in all scenarios for the majority of the datasets. On average, the performance improvement over the supervised baseline methods ranges from about 1% for VPT to about 3% for CoOp with the ResNet-50 backbone. See the Supplementary Material for further analysis of base and combined performance.
@@ 4.5 Results on Class Agnostic Adaptation
Figure 3: Class agnostic adaptation. (a) Comparison between the zero-shot baselines and the proposed class agnostic CA-KDPL variants (highlighted in cyan). The prompt is learned in an unsupervised and class agnostic setting on ImageNet and evaluated on the benchmark datasets used in the domain generalization ($\text{AVG}^{1}$) and cross-dataset ($\text{AVG}^{2}$) evaluations. Average performance improvements are indicated in green, and deterioration in red. (b) Per-dataset accuracy comparison between our unsupervised and class agnostic method (CoOp+CA-KDPL), the supervised baseline CoOp, and the zero-shot student on the cross-dataset benchmark datasets.
@@
Figure 3(a-b) summarizes the main results in the proposed Class-Agnostic (CA) scenario in which even the training class names are unknown at training time. We report the accuracy on the ImageNet dataset (source), as well as the average accuracy in domain generalization and cross-dataset settings. Note that, even without knowing the class names, the performance on the source dataset steadily improves compared to the zero-shot CLIP model. Moreover, the prompts learned via the proposed unsupervised class-agnostic knowledge distillation also exhibit improved average domain generalization ($\text{AVG}^{1}$) and cross-dataset capabilities ($\text{AVG}^{2}$). Figure 3(b) visually depicts how the ResNet-50-based CoOp+CA-KDPL compares with supervised CoOp and zero-shot CLIP student performance in the cross-dataset transfer setting. Notably, CA-KDPL outperforms both baselines despite being unsupervised and class-agnostic during training.
4.6 Ablation Study
@@
(a) KDPL versus pseudolabeling
(b) Sampling strategy
(c) Symmetric/asymmetric KL
Table 4: Ablations on labeling, sampling, and KL-divergence. (a) Comparison of KDPL and a pseudolabeling strategy adapted from [26]. (b) Ablation on a number of classes sampled for class agnostic adaptation. (c) Comparison of forward, reverse, and symmetric KL-divergence used in the loss of Eq. 2.
@@
Pseudolabeling versus KDPL. We compare KDPL and CA-KDPL with a pseudolabeling strategy for exploiting unlabeled samples. Similar to the approach outlined in UPL [26], we generate teacher-derived labels for the training samples. Instead of selecting the top-K confident pseudolabels over the entire training set, for a fair comparison we pseudolabel the fixed, few shot samples in each run. We call this baseline UPL*. Table 4(a) shows that KDPL outperforms UPL*
@@
Sampling Strategies for Class Agnostic Adaptation. An important detail in the class agnostic adaptation scenario is how class names are selected for computing the loss. Existing supervised approaches (e.g. POMP [44]) to handling large numbers of classes suggest simply using the ground truth labels of batch samples and supplementing them with randomly selected ones to reach a predetermined number $K$. We adapted this technique and compare it with ours in Table 4(b). These results show that KDPL outperforms the random selection strategy denoted as POMP*. Notably, the optimal number of classes for domain generalization and cross-dataset adaptation is 1000, suggesting that the optimum closely aligns with the actual number of classes in the training dataset.
@@
Symmetric versus Asymmetric KL. In Eq. 2 we use the symmetric KL-divergence, which is a sum of the forward (from teacher to student) and reverse (from student to teacher) KL-divergences. An asymmetric forward loss may yield benefits when the target distribution is similar to the source distribution (see Table 4(c)). Conversely, a reverse loss may prove better in the opposite scenario. Interestingly, in the class agnostic scenario the optimal strategy is the symmetric KL divergence. This enables learning from both in- and out-of-domain training samples, as demonstrated by the results in Figure 3(a).
5 Conclusion
In this paper we proposed an approach to Knowledge Distillation Prompt Learning (KDPL) which is easily integrated into existing supervised prompt learning methods. Our experiments show that for CoOp, CoCoOp, VPT, and MaPLe adding KDPL: (1) renders them label agnostic by eliminating the need for ground-truth labels for few-shot adaptation; (2) can also render them class agnostic in cases where no knowledge of training class labels is available; and (3) remarkably improves generalization to downstream tasks.
Appendix
A. Overview
This document contains Supplementary Material that provides additional details to the main paper and further experimental analysis. The contents are organized as follows:
  • Additional Implementation Details (Appendix B. Additional Implementation Details): This section provides additional implementation details, including hyperparameters and configuration settings for reproducing all reported results and additional details about the evaluated datasets.
  • Additional Results: Generalization to Unseen Classes (Appendix C. Additional Results: Generalization to Unseen Classes): Here we present additional experimental results for the scenario of generalization from base to unseen classes within the same dataset, including the downstream performance of few-shot base and zero-shot unseen classes.
  • Additional Results: Class Agnostic Scenario (Appendix D. Additional Results: Class Agnostic Scenario): This section provides additional results for the class-agnostic setting, our proposed new setting in which we know neither the labels nor the names of the training classes.
Table 5: Configuration of hyperparameters for the re-computed baselines. Note that CoOp and CoCoOp, being textual prompt learning methods, do not include visual tokens.
Dataset CoOp CoCoOp VPT MaPLe
Batch Size 32 1 4 4
Optimizer SGD SGD SGD SGD
LR 0.02 0.02 0.0025 0.0035
Epochs 50 10 5 5
Text-tokens 4 4 4 4
Visual-tokens 8 2
Vision-depth 12 9
Table 6: Datasets statistics.
Dataset Classes Train Val Test
ImageNet [11] 1,000 1.28M N/A 50,000
Caltech101 [19] 100 4,128 1,649 2,465
OxfordPets [41] 37 2,944 736 3,669
StanfordCars [31] 196 6,509 1,635 8,041
Flowers102 [39] 102 4,093 1,633 2,463
Food101 [5] 101 50,500 20,300 30,300
FGVCAircraft [37] 100 3,334 3,333 3,333
SUN397 [53] 397 15,880 3,970 19,850
DTD [10] 47 2,820 1,128 1,692
EuroSAT [22] 10 13,500 5,400 8,100
UCF101 [50] 101 7,639 1,898 3,783
ImageNetV2 [43] 1,000 N/A N/A 10,000
ImageNet-S [51] 1,000 N/A N/A 50,889
ImageNet-A [24] 200 N/A N/A 7,500
ImageNet-R [23] 200 N/A N/A 30,000
B. Additional Implementation Details
Hyperparameter settings. For all evaluated scenarios we recalculated all the baselines results using the parameter and hyperparameter settings reported in the official publications [65, 64, 28, 29]. In Table 6 we summarize all configuration settings. Notably, all the evaluated baselines use a cosine scheduler with one warm-up epoch with a constant learning rate of 1e-5, and all authors suggest initializing the textual context with “a photo of a” as a better starting point. Note that our variants augmented with Knowledge Distillation Prompt Learning (KDPL) were trained using exactly the same parameters and configuration settings as the baseline methods. Therefore, it is reasonable to expect that further hyperparameter tuning could potentially improve our performance even further.
In addition, all the baselines – and so also our augmented versions – adhere to the same preprocessing pipeline, in which input images are resized to (224, 224) using bicubic interpolation. Data augmentations including random resized crop, random flip, and normalization are employed to ensure robustness across all experiments.
Our implementation is built upon the Dassl framework [^3] https://github.com/KaiyangZhou/Dassl.pytorch, which serves as the foundation for all these methods, to ensure consistency with the baselines and to facilitate easy comparison and evaluation.
We use ResNet-50 and ViT-B/32 as the backbones for all experiments. These backbones are the original publicly available models released by OpenAI [^4] https://github.com/openai/CLIP.
Dataset Details. The detailed statistics of the 11 datasets used in the Cross-dataset Transfer setting, as well as the four variants of ImageNet used in the Cross-domain generalization setting, are shown in Table 6. Note that according to the authors of CoOp, for Caltech101, the “BACKGROUND Google” and “Faces easy” classes are discarded, and for the video dataset UCF101 the middle frame of each video is used as input to the image encoder [65].
C. Additional Results: Generalization to Unseen Classes
@@
Table 7: Base-to-new generalization. Prompts are learned from the base classes (16-shots) and evaluated on the same classes on the test set (Base), zero-shot on the unseen classes of the test set (New), and in terms of harmonic mean (H) between the base and the new accuracy. Here we see how KDPL generalizes better to the new classes, limiting deterioration in performance caused by the unbalanced few-shot training on half the classes of the source domain. Remarkably, we are competitive even on the base classes without requiring training labels.
@@
In the main paper, due to brevity and to maintain focus on the core aspects of our method, we reported only the performance metrics on unseen classes. In this section we provide the performance metrics on base classes as well as the harmonic mean of both base and unseen classes to provide a more comprehensive perspective on our performance in this scenario.
In Table 7, we observe that the baseline methods without KDPL outperform our approach in few-shot transfer learning on base classes. Thanks to the supervision label signal, the baselines can exploit the ground-truth source training classes. Without using KDPL, however, all the baseline methods suffer from forgetting of the unseen classes. In particular, we see that increasing the number of training base-shots overall causes a deterioration in the unseen classes. Certain methods in particular suffer from this phenomenon, typical of an Incremental Learning setting (i.e. in Figure 4 we see that the average performance of MaPLe on unseen classes is consistently below the average zero-shot unseen performance). Applying our method instead exhibits a more favorable trend in zero-shot transfer learning on unseen classes, alleviating the forgetting of unseen classes. Our method is preferable in scenarios where labeled data for base classes are not available, and improvement in base class performance is desired without compromising zero-shot transfer performance on unseen classes.
Figure 4: Generalization to Unseen Classes. The base-to-unseen generalization results of MaPLe+KDPL versus MaPLe are reported. Prompts are learned from the base classes and evaluated on the same base classes and on the unseen classes of the test set. Instead of reporting only the 16-shots performance, we include results for 1-, 2-, 4-, 8-, and 16-shots. Here we see that increasing the number of shots increases the performance of the base class but harms the performance of the unseen classes. However, with our approach we can alleviate this problem, reaching an average 16-shots performance on unseen classes greater than the zero-shot reference, while the average MaPLe performance remains below the average zero-shot reference for all shots.
D. Additional Results: Class Agnostic Scenario
@@
In the main paper we did not include all results for the class agnostic setting due to space limitations, but only reported the average domain generalization performance and the average cross-dataset performance. Here, in Table D. Additional Results: Class Agnostic Scenario we provide all class-agnostic results comparing our proposed method (CA-KDPL) against the baselines. As mentioned in the main paper, even without knowing either the labels or the training class names, our proposed unsupervised method is competitive, and sometimes even outperforms the baselines.

### v2

@@ 4.4 Results on Generalization to Unseen Classes
In Table 3 we give results for the unseen class generalization task. In these experiments each dataset is split into 50% of the classes as a base for training few-shot adaptation, and the remaining 50% as new classes on which zero-shot performance is evaluated. KDPL consistently outperforms the corresponding baseline methods for both backbones, demonstrating improvement in all scenarios for the majority of the datasets. On average, the performance improvement over the supervised baseline methods ranges from about 1% for VPT to about 3% for CoOp with the ResNet-50 backbone. See Section C.1. Generalization to Unseen Classes in the Supplementary Material for further analysis of base and unseen performance.
@@ 4.5 Results on Class Agnostic Adaptation
Figure 3: Class agnostic adaptation. (a) Comparison between the zero-shot baselines and our class agnostic CA-KDPL variants (highlighted in cyan). The prompt is learned in an unsupervised and class agnostic setting on ImageNet and evaluated on the benchmark datasets for domain generalization ($\text{AVG}^{1}$) and cross-dataset ($\text{AVG}^{2}$) evaluations. Average performance improvements are indicated in green, and deterioration in red. (b) Per-dataset accuracy comparison between our unsupervised and class agnostic method (CoOp+CA-KDPL), the supervised baseline CoOp, and the zero-shot student on the cross-dataset benchmark datasets.
@@
Figure 3(a-b) summarizes the main results in the proposed Class Agnostic (CA) scenario in which even the training class names are unknown at training time. We report the accuracy on the ImageNet dataset (source), as well as the average accuracy in domain generalization and cross-dataset settings. Note that, even without knowing the class names, the performance on the source dataset steadily improves compared to the zero-shot CLIP model. Moreover, the prompts learned via the proposed unsupervised class agnostic knowledge distillation also exhibit improved average domain generalization ($\text{AVG}^{1}$) and cross-dataset capabilities ($\text{AVG}^{2}$). Figure 3(b) visually depicts how the ResNet-50-based CoOp+CA-KDPL compares with supervised CoOp and zero-shot CLIP student performance in the cross-dataset transfer setting. Notably, CA-KDPL outperforms both baselines despite being unsupervised and class agnostic during training.
5 Conclusion
In this paper we proposed an approach to Knowledge Distillation Prompt Learning (KDPL) which is easily integrated into existing supervised prompt learning methods. Our experiments show that for CoOp, CoCoOp, VPT, MaPLe and PromptSRC adding KDPL: (1) renders them label agnostic by eliminating the need for ground-truth labels for few-shot adaptation; (2) can also render them class agnostic in cases where no knowledge of training class labels is available; and (3) remarkably improves generalization to downstream tasks.
Acknowledgements
This work was supported by funding from the European Commission Horizon 2020 grant #951911 (AI4Media).
Supplementary Material for:
[…]
Overview
This document contains Supplementary Material that provides additional details and further experimental analysis. The contents are organized as follows:
  • A. Additional Implementation Details Additional Implementation Details: This section provides additional implementation details, including hyperparameters and configuration settings for reproducing all reported results and further details about the datasets used.
  • B. Ablation Study Ablation Study
    • B.1. Pseudolabeling versus KDPL Pseudolabelling versus KDPL: This subsection compares KDPL performances with a famous pseudolabeling prompt learning strategy demonstrating the superiority of KDPL over the baseline approach.
    • B.2. Sampling Strategies for Class Agnostic Adaptation Sampling Strategies for Class Agnostic Adaptation: This part evaluates different sampling strategies for class agnostic adaptation and their performance varying the number of sampled class names.
    • B.3. Symmetric versus Asymmetric KL Symmetric versus Asymmetric KL: This section explores the impact of using symmetric versus asymmetric KL-divergence in the loss function explaining why is more effective for class agnostic scenarios.
    • B.4. Is a smaller Teacher enough? Is a smaller Teacher enough? Here, we investigate whether smaller teachers model can achieve competitive performance compared to the larger ViT-H-14 teacher model used in the main paper.
  • C. Additional Results Additional Results
    • C.1. Generalization to Unseen Classes Generalization to Unseen Classes: Here we present additional experimental results for the generalization to unseen classes within the same dataset scenario, including the downstream performance of few-shot base as well as the harmonic mean of base and unseen classes.
    • C.2. Class Agnostic Scenario Class Agnostic Scenario: This section provides additional results for the class agnostic setting, our proposed new setting in which we know neither the labels nor the names of the training classes.
A. Additional Implementation Details
Table 4: Configuration of hyperparameters for the re-computed baselines. Note that CoOp and CoCoOp, being textual prompt learning methods, do not learn visual tokens.
Dataset CoOp CoCoOp VPT MaPLe PromptSRC
Batch Size 32 1 4 4 4
Optimizer SGD SGD SGD SGD SGD
LR 0.02 0.02 0.0025 0.0035 0.0025
Epochs 50 10 5 5 20
Text-tokens 4 4 4 4 4
Visual-tokens 8 2 4
Vision-depth 12 9 9
Table 5: Datasets statistics.
Dataset Classes Train Val Test
ImageNet [12] 1,000 1.28M N/A 50,000
Caltech101 [20] 100 4,128 1,649 2,465
OxfordPets [43] 37 2,944 736 3,669
StanfordCars [32] 196 6,509 1,635 8,041
Flowers102 [41] 102 4,093 1,633 2,463
Food101 [6] 101 50,500 20,300 30,300
FGVCAircraft [39] 100 3,334 3,333 3,333
SUN397 [56] 397 15,880 3,970 19,850
DTD [11] 47 2,820 1,128 1,692
EuroSAT [23] 10 13,500 5,400 8,100
UCF101 [53] 101 7,639 1,898 3,783
ImageNetV2 [45] 1,000 N/A N/A 10,000
ImageNet-S [54] 1,000 N/A N/A 50,889
ImageNet-A [25] 200 N/A N/A 7,500
ImageNet-R [24] 200 N/A N/A 30,000
Hyperparameter settings. For all evaluated scenarios we recalculated all baseline results using the parameter and hyperparameter settings reported in the original publications [68, 67, 29, 30, 31].
In Table 5 we summarize all configuration settings. Notably, all the evaluated baselines use a cosine scheduler with one warm-up epoch with a constant learning rate of 1e-5, and all authors suggest initializing the textual context with “a photo of a” as a better starting point. Note that our variants augmented with Knowledge Distillation Prompt Learning (KDPL) were trained using exactly the same parameters and configuration settings as the baseline methods. Therefore, it is reasonable to expect that further hyperparameter tuning could potentially improve our performance even further.
In addition, all the baselines – and so also our augmented versions – adhere to the same preprocessing pipeline in which input images are resized to $224\times 224$ pixels using bicubic interpolation. Data augmentations including random resized crop, random flip, and normalization are used to ensure robustness.
To ensure consistency with baselines and to facilitate easy comparison and evaluation, our implementation is built upon the Dassl framework [^3] https://github.com/KaiyangZhou/Dassl.pytorch, which serves as the foundation for all these methods. We use ResNet-50 and ViT-B/32 as the backbones for all experiments. These backbones are the original publicly available models released by OpenAI. [^4] https://github.com/openai/CLIP As mentioned in the main paper, the code is publicly available at https://github.com/miccunifi/KDPL.
Dataset Details.. The detailed statistics of the 11 datasets used in the Cross-dataset transfer setting, as well as the four variants of ImageNet used in the Cross-domain generalization setting, are given in Table 5. Note that according to the authors of CoOp, for Caltech101, the “BACKGROUND Google” and “Faces easy” classes are discarded, and for the video dataset UCF101 the middle frame of each video is used as input to the image encoder [68].
B. Ablation Study
Table 6: Ablations on labeling, sampling, and KL-divergence. (a) Comparison of KDPL and a pseudolabeling strategy adapted from UPL [27]. (b) Ablation on a number of classes sampled for class agnostic adaptation. (c) Comparison of forward, reverse, and symmetric KL-divergence used in the loss (see Eq. 2 of the main paper).
@@ B. Ablation Study
(a) KDPL vs. pseudolabeling (b) Sampling strategy (c) Symmetric/asymmetric KL
@@ B.1. Pseudolabeling versus KDPL
We compare KDPL and CA-KDPL with a pseudolabeling strategy for exploiting unlabeled samples. Similar to the approach outlined in UPL [27], we generate teacher-derived labels for the training samples. Instead of selecting the top-K confident pseudolabels over the entire training set, for a fair comparison we pseudolabel the fixed, few shot samples in each run. We call this baseline UPL*. Table 6(a) shows that KDPL consistently outperforms UPL* on all the evaluated scenarios.
B.2. Sampling Strategies for Class Agnostic Adaptation
@@
An important detail in the class agnostic adaptation scenario is how class names are selected for computing the loss. Existing supervised approaches, such as POMP [46], to handling large numbers of classes suggest simply using the ground truth labels of batch samples and supplementing them with randomly selected ones to reach a predetermined number $K$. We adapted this technique and compare it with ours in Table 6(b). These results show that KDPL outperforms the random selection strategy denoted as POMP*. Notably, the optimal number of classes for domain generalization and cross-dataset adaptation is 1000, suggesting that the optimum closely aligns with the actual number of classes in the training dataset.
B.3. Symmetric versus Asymmetric KL
@@
In Eq. 2 of the main paper we propose to use the symmetric KL-divergence, which is a sum of the forward (from teacher to student) and reverse (from student to teacher) KL-divergences. An asymmetric forward loss may yield benefits when the target distribution is similar to the source distribution (see Table 6(c)). Conversely, a reverse loss may prove better in the opposite scenario. If the Teacher outputs probability zero for a class on a training sample, the forward KL-divergence is zero for that sample even if the Student outputs a very high probability (see Eq. (3), main paper). Adding the reverse KL-divergence prevents the distillation loss from going to zero in such cases (Eq. (2), main paper). Indeed, in the class agnostic scenario the optimal strategy is the symmetric KL divergence. This enables learning from both in- and out-of-domain training samples, as demonstrated by the results in Figure 3 of the main paper.
B.4. Is a smaller Teacher enough?
All the results presented so far involve ViT-H-14 as the teacher. To evaluate if a smaller teacher would be enough we ran additional experiments with ViT-L/14 and ViT-B/16 as Teacher and ResNet-50 as Student (see Table 7). Larger Teachers yield better results, but even more modest ones still maintain excellent performance on domain generalization and cross-dataset transfer. Interestingly, smaller Teachers work better on generalization to unseen classes, which we believe is due to the weaker supervision causing less adaptation and thus maintaining better zero-shot transfer.
Table 7: Average accuracy comparison on the three evaluated scenarios between the supervised baseline CoOp and our unsupervised approach (CoOp + KDPL) using different and smaller teachers.
Target downstream Average
Method Source Domain Generalization Cross-Dataset Transfer Generalization to Unseen Classes
CoOp 62.40 42.05 55.28 60.82
CoOp + KDPL (B/16) 60.73 42.27 56.52 67.57
CoOp + KDPL (L/14) 61.43 42.61 57.10 66.90
CoOp + KDPL (H/14) 62.73 42.90 57.59 63.80
C. Additional Results
@@
Table 8: Base-to-new generalization. Prompts are learned from the base classes (16-shots) and evaluated on the same classes on the test set (Base), zero-shot on the unseen classes of the test set (New), and in terms of harmonic mean (H) between the base and the new accuracy. Average New performance improvements over the baselines are indicated in green. Here we see how KDPL generalizes better to the new classes, limiting deterioration in performance caused by the unbalanced few-shot training on half the classes of the source domain. Remarkably, we are competitive even on the base classes without requiring training labels.
@@
In the following sections we report on additional experiments validating our proposed approach.
C.1. Generalization to Unseen Classes
In the main paper, due to brevity and to maintain focus on the core aspects of our method, we reported only the performance metrics on unseen classes. Here we provide the performance metrics on base classes as well as the harmonic mean of both base and unseen classes to provide a more comprehensive perspective.
In Table 8, we observe that the baseline methods without KDPL outperform our approach in few-shot transfer learning on base classes. Thanks to the supervision label signal, the baselines can exploit the ground-truth source training classes. Without using KDPL, however, all the baseline methods suffer from forgetting the unseen classes. In particular, we see that training on base-shots causes a decrease in performance on the unseen classes. Certain methods in particular suffer from this phenomenon, typical of an Incremental Learning setting (e.g. in Figure 4 we see that the average performance of MaPLe on unseen classes is consistently below the average zero-shot unseen performance). Applying our method instead exhibits a more favorable trend in zero-shot transfer learning on unseen classes, alleviating the forgetting of unseen classes. Our method is preferable in scenarios where labeled data for base classes are not available, and improvement in base class performance is desired without compromising zero-shot transfer performance on unseen classes.
Figure 4: Generalization to Unseen Classes. The base-to-unseen generalization results of MaPLe+KDPL versus MaPLe are reported. Prompts are learned from the base classes and evaluated on the same base classes and on the unseen classes of the test set. Instead of reporting only the 16-shot performance, we include results for 1-, 2-, 4-, 8-, and 16-shots. Here we see that increasing the number of shots increases the performance of the base class but harms the performance on unseen classes. However, with our approach we can alleviate this problem, reaching an average 16-shots performance on unseen classes greater than the zero-shot reference, while the average MaPLe performance remains below the average zero-shot reference for all shots.
C.2. Class Agnostic Scenario
@@
In the main paper we did not include all results for the class agnostic setting due to space limitations, but only reported the average domain generalization performance and the average cross-dataset performance. Here, in Table 9 we provide all class agnostic results comparing our proposed method (CA-KDPL) against the baselines. As mentioned in the main paper, even without knowing either the labels or the training class names, our proposed unsupervised method is competitive, and sometimes even outperforms the baselines.